Boudreaux is hunting quail with his shotgun on a platform 16 feet above the ground. The quail fly at a height between 1160ft and 1200ft above the ground. The height (in feet) of the shots above the ground at any time (t in seconds) is given by: H(t)=-16t^2+512t+16
Is Boudreaux's shotgun powerful enough to get him a quail?

Answer by stanbon(57219)

Boudreaux is hunting quail with his shotgun on a platform 16 feet above the ground. The quail fly at a height between 1160ft and 1200ft above the ground. The height (in feet) of the shots above the ground at any time (t in seconds) is given by: H(t)=-16t^2+512t+16
Is Boudreaux's shotgun powerful enough to get him a quail?
------------------------------
The maximum height for the gun occurs when t = -b/2a = -512/(2*-16) = 16
------------
The height at t = 16 is -16(16)^2+512*16+16 = 4112 ft
--------------
The gun has plenty of power; certainly enough to reach between 1160 and 1200'
========================================================
Cheers,
Stan H.